you want a one million dollar nest egg when you retire. At present, you have $30,000 saved and would like to retire in 20 years. to reach your goal, you have arbitrarily chosen to set aside $300 a month towards your goal. if you stay on track with your current plan, assuming an average annual return of 6%, how much will you have saved by retirement?
Answer #1

FV of Savings = Amount Saved * (1 + r)n = $30,000 * (1 + 0.06)20 = $30,000 * 3.2071 = $96,214.06

FV of Annuity = Monthly Payment / [{(1 + r)n - 1} / r]

= $300 * [{(1 + 0.06/12)(20*12) - 1} / (0.06/12)]

= $300 * [2.3102 / 0.005]

= $300 * 462.0409

= $138,612.27

Amount saved by retirement = FV of Savings + FV of Annuity

= $96,214.06 + $138,612.27 = $234,826.33